### v2

# DCG-SQL: Enhancing In-Context Learning for Text-to-SQL with Deep Contextual Schema Link Graph

Abstract
Text-to-SQL, which translates a natural language question into an SQL query, has advanced with in-context learning of Large Language Models (LLMs). However, existing methods show little improvement in performance compared to randomly chosen demonstrations, and significant performance drops when smaller LLMs (e.g., Llama 3.1-8B) are used. This indicates that these methods heavily rely on the intrinsic capabilities of hyper-scaled LLMs, rather than effectively retrieving useful demonstrations. In this paper, we propose a novel approach for effectively retrieving demonstrations and generating SQL queries. We construct a Deep Contextual Schema Link Graph, which contains key information and semantic relationship between a question and its database schema items. This graph-based structure enables effective representation of Text-to-SQL samples and retrieval of useful demonstrations for in-context learning. Experimental results on the Spider benchmark demonstrate the effectiveness of our approach, showing consistent improvements in SQL generation performance and efficiency across both hyper-scaled LLMs and small LLMs. The code is available at https://github.com/jjklle/DCG-SQL.
DCG-SQL: Enhancing In-Context Learning for Text-to-SQL
with Deep Contextual Schema Link Graph
Jihyung Lee1thanks: Equal contribution, Jin-Seop Lee [^1] footnotemark: 1, Jaehoon Lee1, YunSeok Choi2thanks: Corresponding author, Jee-Hyong Lee322footnotemark: 2 1Department of Artificial Intelligence, Sungkyunkwan University 2Department of Immersive Media Engineering, Sungkyunkwan University 3Department of Computer Science and Engineering, Sungkyunkwan University {jjklle, wlstjq0602, hoon1223, ys.choi, john}@skku.edu
1 Introduction
Text-to-SQL aims to generate SQL queries given a pair of a natural language question and a database schema. Traditional studies focused on how to train models that took a question and a schema as input to generate the target SQL Wang et al. (2020); Qi et al. (2022); Li et al. (2023b). However, they required training an encoder-decoder model with billions of parameters, leading to high training costs and limited generalization. Recently, with the advance of Large Language Models (LLMs), in-context learning with LLMs has been applied to Text-to-SQL tasks due to its efficiency and strong generalization ability across various tasks Rajkumar et al. (2022); Chang and Fosler-Lussier (2023); Pourreza and Rafiei (2023).
Methods Retrieval GPT-4 Llama 3.1-8B
ACT-SQL 83.9 75.0
83.2 75.6
DAIL-SQL 82.1 69.9
80.8 69.3
Table 1: Execution Accuracy (%) for Spider Dataset. All experiments are conducted with 4-shot setting. When ✗ is applied, the demonstrations are randomly choosen.
To maximize the capability of in-context learning of LLMs, it is necessary to provide relevant and useful demonstrations to generate correct SQL query Luo et al. (2024). Most existing approaches represent each sample using a text embedding derived solely from the question, and select demonstrations based on the similarity of the question embeddings Lee et al. (2025); Guo et al. (2024); Zhang et al. (2023a); Gao et al. (2024). However, as shown in Table 1, their performance shows no difference from that of randomly chosen demonstrations. This result implies that existing methods failed to maximize the capabilities of in-context learning due to their limitation in selecting appropriate demonstrations. This becomes more apparent when they are applied to smaller LLMs which have weaker in-context learning capabilities compared to hyper-scale LLMs Wei et al. (2022). Since smaller LLMs rely more heavily on demonstrations, performance significantly drops, sometimes even falling below the performance with randomly selected demonstrations.
This is because they overlook DB schema when representing Text-to-SQL samples. Since DB schema determines the structure and semantics of SQL queries, it must be incorporated into the representation. In Text-to-SQL tasks, understanding a question inherently relies on the database schema. Therefore, when retrieving relevant demonstrations, both the question and the schema should be jointly considered and represented in a unified manner. encoded through a unified representation.
In this paper, we propose a novel Text-to-SQL method for in-context learning with Deep Contextual Schema Link Graph-based Retrieval, as we called DCG-SQL. We develop a novel graph-based joint representation of both the question and the schema to retrieve demonstrations. However, developing a joint representation that effectively captures the contextual relationship between the question and the schema presents two challenges.
First, datasets are not usually available to learn how schema items are relevant to the question, a process commonly known as schema linking Guo et al. (2019); Wang et al. (2020). As a result, some previous works Qi et al. (2022); Li et al. (2023b); Zhang et al. (2023b) have attempted to establish links through simple word matching, but this approach is often too simplistic and fails to accurately capture the relationship between questions and schema items. A novel method is needed to effectively model the contextual relationships necessary for Text-to-SQL samples without annotations.
The second challenge is not all schema items in a database are necessary for effectively representing the given Text-to-SQL sample. If all schema elements are considered, irrelevant schema items could hinder the effective representation. If irrelevant schema items are included in representations, question tokens could be connected to unrelated schema items. They also make it difficult to distinguish between samples with different questions that share the same DB schema. Therefore, when representing Text-to-SQL samples for demonstration retrieval, it is also essential to exclude irrelevant schema items.
To identify only relevant schema items to the question in a sample, we train a transformer based encoder with classification heads. Since each training sample contains a question, a schema and a target SQL, we can train the classifier with the question and the schema as input, and the SQL to indicate relevant schema items. When training the encoder, we add special tokens, such as [TAB] and [COL], to generate appropriate representations even when their names are composed of several words. These special tokens help to effectively capture their semantics. Next, we generate a graph with the selected schema items and the words in the question. We connect items and words using attention scores between question tokens and schema item tokens. Since the attention mechanism enables the model to assign relevance between tokens Galassi et al. (2020), we use the attention scores to predict the schema linking. This allows the model to capture the contextual relationships without any explicit linking dataset. With these, we can effectively achieve a joint representation of each text-to-SQL sample as a graph. By comparing the representations of the graphs, we can retrieve demonstrations. With the demonstrations, our method fully leverages in-context learning capability and generate the target SQL correctly.
To evaluate our proposed method, we conduct experiments on various benchmark datasets, including Spider, Spider-DK, Spider-Realistic, and Spider-Syn Yu et al. (2018b); Gan et al. (2021b); Deng et al. (2021); Gan et al. (2021a). Moreover, we experiment with various hyper-scaled and small LLMs, including GPT-4, GPT-3.5-Turbo, DeepSeek-Coder-33B, Llama 3.1-8B, DeepSeek-Coder-6.7B, and Llama 3.2-3B OpenAI et al. (2024); DeepSeek-AI et al. (2024); Dubey et al. (2024). Our method consistently achieves performance improvements not only with hyper-scaled LLMs but also with small LLMs.
2 Related Work
Figure 1: Overview of our proposed method for Text-to-SQL. The method consists of three main steps: (1) Deep Contextual Schema Link Graph Construction, which prunes irrelevant schema items and builds a schema link graph that captures the contextual relationship between the question tokens and the schema items; (2) Graph-based Demonstration Retrieval, where graph embeddings are used to retrieve demonstrations relevant to the given query; and (3) SQL Generation using retrieved demonstrations to generate the final SQL query.
2.1 In-context Learning for Text-to-SQL
With the advancement of large language models, there has been a growing focus on leveraging them to solve Text-to-SQL tasks Pourreza and Rafiei (2023); Chang and Fosler-Lussier (2023); Guo et al. (2024); Zhang et al. (2023a); Lee et al. (2025); Gao et al. (2024); Nan et al. (2023); Li et al. (2024); Shen et al. (2024). Some approaches decompose the Text-to-SQL task into multiple steps and employ GPT-4 at each step Pourreza and Rafiei (2023); Li et al. (2024); Lee et al. (2025). However, this repeated use of GPT-4, a hyper-scaled LLM, results in significantly high inference costs. Also, since they heavily depend on the hyper-scaled LLM’s capabilities, their performance significantly drops with smaller LLMs, as limited capability of small LLM at each step leads to error accumulation throughout the process.
On the other hand, others Zhang et al. (2023a); Guo et al. (2024); Nan et al. (2023); Gao et al. (2024); Shen et al. (2024); Li et al. (2024); Lee et al. (2025) tried to leverage in-context learning capability of LLMs by retrieving relevant samples and providing them as demonstrations. Zhang et al. (2023a); Guo et al. (2024); Li et al. (2024); Lee et al. (2025) encoded the question into a text embedding using a pre-trained text encoder and retrieved demonstrations based on the embedding similarity. However, they did not consider the DB schema, which is also crucial in Text-to-SQL. In contrast, Gao et al. (2024); Nan et al. (2023); Shen et al. (2024) tried to reflect DB schema information indirectly by approximating target SQL of user input and retrieving similar examples through similarity comparison of the approximated SQL. However, they still have the limitation of requiring costly training and relying on the approximator which has limited capacity for generating SQL. As a result, inaccurate DB schema information can be represented when the approximator generates wrong SQL. Since they did not effectively consider the DB schema during retrieval, they failed to provide useful demonstrations for LLMs’ in-context learning.
2.2 Graph representation for text-to-SQL samples
Before the advent of LLMs, Text-to-SQL research learned schema linking information, which explicitly associates question tokens with relevant database schema elements, to generate target SQL using encoder-decoder models. This helps the model interpret the question and recognize how schema components are organized within the database, leading to improve SQL generation accuracy Lei et al. (2020).
To provide schema linking information to the model, Wang et al. (2020); Qi et al. (2022); Zhang et al. (2023b); Li et al. (2023b) represented the input in a graph format, reflecting the relationships between question tokens and database schema elements. Due to the lack of datasets that explicitly annotate schema linking, these methods were based on simple textual matching, comparing only the spelling between question tokens and schema items. However, such simple linking often fails to capture deeper contextual relationship when question token and schema items are expressed in different forms.
Furthermore, their graph representations contain all schema items, even those unrelated to the question, which can lead to incorrect schema linking. Moreover, retrieval becomes challenging when all schema items are included, as each graph is mostly composed of DB schema elements, making samples that share the same DB schema indistinguishable. Therefore, to retrieve useful demonstrations, a more effective graph representation is needed—one that captures deeper contextual relationships while including only the relevant schema elements.
Figure 2: Deep Contextual Schema Link Graph Construction.
3 Proposed Method
In this section, we introduce a Text-to-SQL method with Deep Contextual Schema Link Graph-based Retrieval, DCG-SQL. Figure 1 illustrates an overview of our method, consisting of three main processes: Deep Contextual Schema Link Graph Construction, Graph-based Demonstration Retrieval, and SQL Generation.
First, we construct a deep contextual schema link graph, which captures the contextual relationships between question and relevant schema items. Then, to retrieve demonstrations with our graph representation, we train a graph encoder with self-supervised learning. This ensures that samples closer to the anchor are more useful in generating the target SQL, enabling the effective retrieval of useful demonstrations. Finally, with the selected demonstrations, our method can fully leverage in-context learning ability, and generate the target SQL more correctly.
3.1 Deep Contextual Schema Link Graph Construction
We propose a deep contextual schema link graph construction method that excludes irrelevant schema items and extracts only the relevant ones. Our method also links the question tokens with the schema items based on their contextual relation.
Schema pruning
DB schema consists of multiple tables and columns, but only a small subset is relevant for SQL generation. If irrelevant DB schema items are included in the graph, they could hinder extracting representation and retrieving useful demonstrations. To focus on the relevant components, we introduce a schema pruning method that identifies the relevant schema items necessary for generating the SQL.
For schema pruning, we utilize a cross-encoder model to determine the relevance of tables and columns Li et al. (2023a). The input to the cross-encoder is a flattened text sequence that combines the question with schema items (either a table or a column). Since schema items, such as the column name ‘head id’, may be tokenized into multiple tokens, they need to be represented as a single token. We prepend special tokens [TAB] and [COL] for representing tables and columns, respectively. The special token embeddings for schema items are initialized by applying mean pooling over the token embeddings of each word in the schema item. Unlike existing pruning approach Li et al. (2023a), these special tokens enable us to leverage the attention scores between question and schema items.
Let $Q$ denote the tokens of the question, and $T$ and $C$ denote the names of the table and column. The input sequence $X$ is represented as:
$X$ =$=\{Q\,\textnormal{[TAB]}\,T_{1}\,\textnormal{[COL]}\,C_{11},\ldots,$
[…]
where $T_{i}$ denotes the $i$-th table in the DB schema, and $C_{ij}$ refers to the $j$-th column of the table $T_{i}$. The cross-encoder model is trained as binary classification for each special token to predict the relevant tables and columns for a given question, as shown in Figure 2.
Schema linking
Identifying schema items contextually relevant to the question is crucial for understanding the question and generating SQL outputs Lei et al. (2020); Li et al. (2023b); Wang et al. (2020). However, due to the absence of ground truth dataset for such relation, existing approaches rely on simple text correspondence Li et al. (2023b); Zhang et al. (2023b); Qi et al. (2022); Wang et al. (2020). It compares only the spelling between question tokens and schema items, often failing to capture links when schema items are phrased differently (e.g., linking the question tokens "older" and "56" to the column "age" in Figure 2).
To overcome this issue and capture contextual similarities, we leverage the attention scores from the schema pruning model. Since the pruning model is trained to identify relevant schema items for a given question, its attention scores implicitly reflect the relevance between them. If the schema pruning model classifies a schema item as relevant to SQL generation, we obtain the attention scores between its concatenated special token and the question tokens.
When the attention score exceeds the threshold $\tau$, a link is connected between the schema item and the question token, which we refer to as an ‘attention-match’. This approach enables schema pruning and linking simultaneously without ground truth dataset or additional training. Also, we link tokens that have high syntactic correlation for edges between question and all schema items within each table for edges between schema items, as following existing approaches Qi et al. (2022); Zhang et al. (2023b); Li et al. (2023b); Wang et al. (2020).
Finally, we construct a deep contextual schema linking graph $G$, where question and schema are jointly represented. The graph consists of nodes, which correspond to question tokens and schema items, and edges that capture the contextual relationships between them.
3.2 Graph-based Demonstration Retrieval
To retrieve useful demonstrations based on the deep contextual schema link graph, we propose a graph-based retrieval method that jointly consider the question and the DB schema. To achieve this, we adopt a graph encoder to effectively represent each sample’s graph into an embedding vector Wang et al. (2020); Qi et al. (2022).
To effectively train the graph encoder, we utilize contrastive learning that clusters similar graphs closely in the vector space while distancing dissimilar ones farther apart. We define positive demonstration samples as those that are useful for generating the SQL query of a given anchor. To achieve this, for the given anchor sample $x=(G_{x},SQL_{x})$, we extract likelihoods of the $SQL_{x}$ when other samples $S=\{s_{i}=(G_{i},SQL_{i})\mid i=1,2,\dots\}$ are provided as demonstrations using Llama 3.1-8B model Dubey et al. (2024). The likelihood is as follows:
$score(x,s_{i})=P_{\text{sLLM}}(SQL_{x}\mid G_{x};s_{i})$ (1)
If the likelihood of $SQL_{x}$ is high, it indicates that $s_{i}$ is useful for generating the target SQL query.
To identify positive and negative samples, we rank $s_{i}$’s based on their likelihood of generating the anchor’s target SQL, treating high-ranked ones as positives and low-ranked ones as negatives. Since the retrieval model is trained to increase similarity with high-ranked demonstrations and decrease similarity with low-ranked ones, we can retrieve useful demonstrations from the retrieval pool through similarity of the graph embeddings Karpukhin et al. (2020); Li et al. (2023c).
However, computing the likelihood of $SQL_{x}$ over the entire training dataset as demonstrations is computationally expensive. To address this, we construct a reduced candidate set for each anchor by selecting 100 structurally similar SQLs, using a tree edit distance computed over abstract syntax trees Yu et al. (2018a). Unlike Shen et al. (2024), which selects demonstrations based on AST similarity, we further re-rank the candidates by their usefulness for generating the anchor SQL. This enables us to identify samples that appear structurally similar but are unhelpful as demonstrations—performing hard negative mining.
During inference, we compare the embedding of user input graph $G_{user}$ generated by the encoder, with the embeddings of the demonstration pool by computing the similarity. We select top-k samples with the highest similarity as demonstrations.
3.3 SQL generation with In-context Learning
With the retrieved demonstrations, LLM can generate target SQL query correctly. In addition, we utilized automated-CoT prompting method to guide the SQL generation more effectively Zhang et al. (2022, 2023a). The retrieved demonstrations are categorized into types such as simple, join, and nested queries. Each type uses a predefined CoT template, dynamically adapted per demonstration to guide consistent and structured SQL generation. Details are provided in Appendix C.2.
Method Easy Medium Hard Extra All
EX EM EX EM EX EM EX EM EX EM
Finetuned Model
Graphix-T5 92.3 91.9 86.3 82.3 73.6 65.5 57.2 53.0 80.9 77.1
RESD-SQL - - - - - - - - 84.1 80.1
GPT-4
DIN-SQL 92.3 82.7 87.4 65.5 76.4 42.0 62.7 30.7 82.8 60.1
ACT-SQL 92.3 85.9 89.3 59.1 80.2 53.1 61.5 18.9 83.9 57.9
DAIL-SQL 91.1 90.3 88.6 76.0 75.9 60.3 62.0 50.0 82.8 72.6
ASTRES - - - - - - - - 86.6 77.3
DCG-SQL (Ours) 95.6 92.3 90.6 83.0 82.8 70.7 72.3 61.4 87.5 79.7
GPT-3.5-Turbo
DIN-SQL 85.1 56.5 81.8 51.6 62.1 28.2 51.2 7.8 74.4 41.8
ACT-SQL 92.7 87.9 86.1 52.5 70.1 41.4 56.6 13.3 80.3 52.8
DAIL-SQL 90.3 89.9 82.7 57.0 73.0 52.3 56.0 18.7 78.0 64.9
ASTRES - - - - - - - - 83.0 68.8
DCG-SQL (Ours) 91.9 88.3 89.2 78.0 74.7 56.3 63.3 39.8 83.3 70.7
Deepseek-Coder-33B-Instruct
DIN-SQL 91.9 81.9 82.1 62.3 58.0 31.0 50.0 17.5 75.0 54.5
ACT-SQL 84.3 78.6 85.0 57.4 64.4 44.8 50.6 25.3 75.8 55.2
DAIL-SQL 90.3 87.9 86.5 73.8 73.0 62.6 57.2 45.2 80.5 70.7
ASTRES - - - - - - - - 83.4 64.7
DCG-SQL (Ours) 94.0 90.7 92.6 80.0 75.3 63.8 63.3 53.0 85.3 75.5
Llama 3.1-8B-Instruct
DIN-SQL 84.7 79.4 72.2 54.9 60.9 36.8 35.5 19.9 67.4 52.1
ACT-SQL 89.5 85.9 81.8 63.2 60.9 42.0 50.0 31.9 75.0 60.1
DAIL-SQL 81.9 67.7 75.3 57.4 60.9 37.9 54.8 46.4 71.2 54.8
DCG-SQL (Ours) 94.8 93.1 87.0 78.7 77.0 57.5 55.4 43.4 82.1 72.9
Deepseek-Coder-6.7B-Instruct
DIN-SQL 86.7 70.6 78.7 57.6 59.2 29.3 49.4 21.1 72.6 50.1
ACT-SQL 88.7 82.7 81.4 57.4 66.7 46.6 41.0 20.5 74.2 55.7
DAIL-SQL 85.5 71.0 79.1 63.0 64.9 47.7 48.8 41.6 73.4 58.9
DCG-SQL (Ours) 91.9 89.9 86.8 78.0 74.7 56.3 56.6 48.8 81.1 72.5
Llama 3.2-3B-Instruct
DIN-SQL 69.4 60.1 47.5 35.2 31.6 17.8 14.5 3.6 44.8 33.2
ACT-SQL 85.9 77.8 71.7 56.3 53.4 37.4 33.7 24.1 66.0 53.1
DAIL-SQL 68.1 50.8 63.7 49.6 44.8 25.9 46.4 42.8 58.8 44.8
DCG-SQL (Ours) 91.9 87.5 79.8 71.7 60.9 48.9 50.6 41.0 74.7 66.7
Table 2: Experimental Results on Spider Dataset with various LLMs.
4 Experiments
4.1 Implementation Details
All experiments are conducted on an A100 80GB GPU. Our method requires only small size of models for efficiency. For schema link graph construction, we train the schema pruning model based on RoBERTa-large, which has approximately 355 million parameters Liu et al. (2019). For schema linking, we set $\tau_{tab}$ and $\tau_{col}$ to 0.66 and 0.43, respectively. For graph-based demonstration retrieval, We train a graph encoder model based on a bi-encoder architecture Karpukhin et al. (2020), where each encoder follows the Relation-Aware Transformer Shaw et al. (2018) structure initialized with BERT-base Devlin et al. (2019), totaling approximately 220 million parameters.
For a fair comparison, we reproduce experimental results or follow existing studies. For SQL generation, we use various LLMs, including GPT-4, GPT-3.5-Turbo, DeepSeek-Coder-33B-Instruct, Llama 3.1-8B-Instruct, DeepSeek-Coder-6.7B-Instruct, and Llama 3.2-3B-Instruct OpenAI et al. (2024); DeepSeek-AI et al. (2024); Dubey et al. (2024). In all experiments, the generation temperature is set to 0. We set 5 demonstrations for the few-shot setting and our input prompt is introduced in Appendix C.1. Training the pruning model took approximately 3 hours, while the graph encoder model was trained for 8 hours. We used spaCy Honnibal and Montani (2017) and NLTK Bird and Loper (2004) libraries for the edges between question tokens.
4.2 Dataset Details
To verify the effectiveness of our method, we employ the SPIDER dataset Yu et al. (2018b), large-scale benchmark designed for the text-to-SQL task. Spider spans 200 distinct databases across 138 domains and is widely regarded for its diverse and cross-domain nature. The dataset comprises 7,000 training examples and 1,034 development examples, providing a robust basis for model assessment. The dataset categorizes SQL queries based on their complexity into four difficulty levels: easy, medium, hard, extra hard. Our evaluation experiments are conducted on the development set, using the training set as the demonstration pool for retrieval. Moreover, using the same demonstration pool, we evaluate on Spider-DK Gan et al. (2021b), Spider-Realistic Deng et al. (2021), and Spider-Syn Gan et al. (2021a), which are derived from Spider. They incorporate domain knowledge, remove explicitly mentioned column names, and replace schema-related words with synonyms, respectively.
Method DK Real Syn
EX EM EX EM EX EM
GPT-4
ACT-SQL 72.0 47.7 81.3 55.7 72.1 48.5
DAIL-SQL - - 76.0 - - -
ASTRES 72.3 59.1 80.9 66.1 74.4 61.3
DCG-SQL 71.6 66.5 81.9 73.6 78.7 68.6
GPT-3.5-Turbo
DIN-SQL 64.3 36.6 65.2 41.1 65.3 34.7
ACT-SQL 68.2 45.4 76.4 48.6 70.4 44.4
DAIL-SQL - - 67.9 - - -
ASTRES 68.8 49.3 78.0 60.8 66.9 51.2
DCG-SQL 69.0 59.4 79.0 62.8 74.2 61.6
Deepseek-Coder-33B-Instruct
ACT-SQL 69.5 47.5 76.4 51.8 69.1 49.3
ASTRES 70.5 46.4 77.4 59.3 68.7 49.5
DCG-SQL 68.4 62.8 79.1 71.1 73.6 62.6
Llama-3.1-8B-Instruct
DIN-SQL 56.3 44.1 59.6 36.4 57.2 43.5
ACT-SQL 63.6 48.0 74.4 60.2 66.6 53.7
DCG-SQL 66.5 60.2 75.8 67.1 70.1 59.7
Deepseek-Coder-6.7B-Instruct
DIN-SQL 54.1 16.2 53.5 29.3 53.7 25.4
ACT-SQL 63.2 44.9 69.1 53.5 63.2 44.6
DCG-SQL 65.0 55.5 74.8 68.5 70.5 59.9
Llama-3.2-3B-Instruct
DIN-SQL 35.7 27.5 38.6 29.1 39.0 27.2
ACT-SQL 54.4 41.7 62.0 47.0 57.4 44.4
DCG-SQL 59.1 52.5 68.5 59.6 60.9 51.9
Table 3: Experimental Results on Spider-DK, Spider-Realistic, and Spider-Syn with various LLMs.
4.3 Baselines & Evaluation Metrics
We compared our method with recent state-of-the-art in-context learning baselines for Text-to-SQL Pourreza and Rafiei (2023); Zhang et al. (2023a); Gao et al. (2024); Shen et al. (2024).
We use two evaluation metrics, Execution Accuracy (EX) and Exact-set-match Accuracy (EM). Execution accuracy measures whether the predicted SQL query yields the same result as the ground truth query when executed. Exact-set-match accuracy treats each clause as a set and compares the predicted clause to the ground truth clause.
4.4 Experimental Results
Table 2 presents the Text-to-SQL performance on the Spider dataset using various hyper-scaled LLMs and small LLMs. Our proposed method consistently achieves higher execution accuracy and exact match accuracy compared to existing approaches. In particular, our method achieves notable improvements for sLLMs. For execution accuracy (EX), it increases by 9.5% with Llama 3.1-8B and 9.3% with DeepSeek-Coder-6.7B. Notably, there is a substantial gap in exact match accuracy (EM), increasing by 21.3% with Llama 3.1-8B and 23.1% with DeepSeek-Coder-6.7B. This suggests that our demonstrations provide the precise patterns needed for the test cases, as in-context learning relies on pattern following Brown et al. (2020); Luo et al. (2024). These results indicate that our approach, which leverages deep contextual schema link graph-based retrieval for SQL generation, effectively utilizes the in-context learning capabilities, leading to outperforming existing methods in text-to-SQL task.
Additionally, to demonstrate the effectiveness of our approach in more challenging scenarios, we conduct experiments on Spider-DK, Spider-Realistic, and Spider-Syn for SQL generation, as shown in Table 3. Our method consistently demonstrates strong performance across various LLM scenarios and datasets. These findings highlight the effectiveness of our method across different datasets, showing strong performance not only with hyper-scaled LLMs but also with small LLMs.
4.5 Ablation Studies
Retrieval Spider Real Syn
EX EM EX EM EX EM
GPT-3.5-Turbo
Random 78.8 47.0 74.6 42.2 60.0 38.4
DCG-SQL 83.3 70.7 79.0 62.8 74.2 61.6
Llama-3.1-8B-Instruct
Random 73.0 48.5 69.7 45.5 61.6 39.7
DCG-SQL 82.1 72.9 75.8 67.1 70.1 59.7
Table 4: Effectiveness of our retrieved demonstrations.
Retrieval Case EX (%) EM (%)
(1) 77.0 66.2
(2) 77.1 65.6
(3) 76.2 66.5
(4) 76.1 59.6
Ours 82.1 72.9
Table 5: Effectiveness of our deep contextual schema link graph-based retrieval. The experiments were conducted on Spider with Llama-3.1-8B-Instruct.
DB Schema Type Precision Recall F1 Score
Table 96.2 97.4 96.8
Column 90.7 95.5 93.1
Table 6: Schema pruning performance.
Method EX (%) EM (%)
DAIL-SQL 71.2 54.8
DAIL-SQL w/ Our Retrieval 75.6 61.3
Table 7: Transferability of our proposed modules. The experiments were conducted on Spider with Llama-8B.
Method Average Tree Edit Distance
ACT-SQL 27.17
DAIL-SQL 22.38
DCG-SQL 13.82
Table 8: Average Tree Edit Distance between demonstration and target SQL query on Spider Dataset
Effectiveness of our retrieval method.
Table 4 presents the results when demonstrations are randomly chosen. Existing approaches Zhang et al. (2023a); Gao et al. (2024) failed to maximize in-context learning capabilities due to their ineffective retrieval approaches as shown in Table 1, whereas our method consistently achieves performance improvements across various datasets and LLMs.
To demonstrate the effectiveness of our retrieval method, we compare it with various retrieval approaches, as shown in Table 5. For demonstration retrieval, case (1) is based on the text embedding of question and (2) is based on the masked question text embedding Guo et al. (2024). Case (3) is based on both question and DB schema, but uses text embeddings. In case (4), it is based on the existing graph representation Li et al. (2023b). Cases (1)–(3) use BERT-base as the text encoder, whereas case (4) employs the same graph encoder as our method. All (1)–(4) are trained using the same self-supervised learning approach described in Section 3.2
As shown in cases (1) and (2), demonstration retrieval based only on question does not achieve good performance. Also, retrieval methods that consider both question and DB schema information without their contextual relationships, show poor performance. In contrast, our method captures contextual relationships in the joint representation of Text-to-SQL samples, leading to better demonstration retrieval.
Effectiveness of our DB schema pruning.
We demonstrate the effectiveness of our pruning method as shown in Table 6. For tables, our method achieves a precision of 96.2%, recall of 97.4%, and an F1 score of 96.8%. For columns, our method shows a precision of 90.7%, recall of 95.5%, and an F1 score of 93.1%. These results show that our method effectively identifies relevant schema items and exclude irrelevant links.
Transferability of our proposed modules.
To demonstrate the transferability of our modules, we applied our retrieval to the DAIL-SQL Gao et al. (2024). All the experimental results in Table 7 are obtained using the same setting as DAIL-SQL. When we use the demonstrations sets retrieved by our method, the performance is 75.6% in EX and 61.3% in EM, respectively. This corresponds to improvements of 6.2% and 11.9% over the original DAIL-SQL. These results demonstrate that our retrieval method can be effectively applied to various prompt templates and Text-to-SQL methods.
Analysis of retrieved demonstrations.
To analyze the quality of retrieved demonstrations, we calculate the average Tree Edit Distance (TED) Yu et al. (2018a), which measures the similarity between the target SQL of user input and the SQLs of the retrieved demonstrations. If TED is small, it indicates that the SQLs of retrieved demonstrations are similar to the target SQL of user input. As shown in the Table 8, our method achieves the lowest TED compared to existing approaches. Since in-context learning of LLMs tends to follow the patterns in the demonstrations Brown et al. (2020); Luo et al. (2024), the result supports that our proposed method effectively retrieves useful demonstrations for generating correct SQL.
Method EX (%) Retrieval Params. LLM Calls Retrieval Latency
DIN-SQL 67.4 - 4 -
ACT-SQL 75.0 1M 1 0.7
DAIL-SQL 71.2 3B 1 9.1
DCG-SQL 82.1 0.5B 1 1.3
Table 9: Comparison of Efficiency. Experimental results are measured on Spider with Llama-3.1-8B-Instruct.
Efficiency of our retrieval method.
To demonstrate the efficiency of our method, we evaluate the inference latency on the Spider dataset, as shown in Table 9. Retrieval latency refers to the time required on average to extract embeddings from test cases. DIN-SQL Pourreza and Rafiei (2023), which performs four separate LLM inferences without retrieval, incurs four times the LLM latency compared to other methods. Meanwhile, DAIL-SQL Gao et al. (2024) employs a retrieval model with 3B parameters Li et al. (2023b), resulting in latency up to 9.1 seconds. Although ACT-SQL Zhang et al. (2023a) shows relatively low inference latency, its execution accuracy remains unsatisfactory. In contrast, our method requires only a single LLM inference and employs lightweight retrieval modules with just 0.5B trainable parameters, including the schema pruning model and the graph encoder. This design enables more efficient retrieval while maintaining high execution accuracy. These results clearly demonstrate the effectiveness and efficiency of our approach.
5 Conclusion
This paper presents a novel approach for retrieving demonstrations and generating SQL queries, effective across LLMs of various sizes. Our method leverages a deep contextual schema link graph to enhance demonstration retrieval and SQL generation. It achieved strong performance across hyper-scaled and small LLMs.
Limitations
Under the same LLM models, our method outperforms the existing SOTA. Our method has shown highly effective in improving performance with small LLMs, however, there is still a performance gap between the use of hyper-scaled LLMs and small LLMs. Future work should focus on addressing this performance gap effectively.
Acknowledgments
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.IITP-2025-RS-2025-00437633, Information Technology Research Center, 20%), (No.RS-2019-II190421, AI Graduate School Support Program(Sungkyunkwan University), 20%), (No.RS-2025-02218768, Accelerated Insight Reasoning via Continual Learning, 20%). This work was also supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No.RS-2025-00521391, 20%) and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea)&Gwangju Metropolitan City (20%).
Appendix A Generalization Performance on BIRD dataset
To verify the generalization ability of our method, we conduct experiments on the BIRD dataset using the same hyperparameters and evaluate its performance on the BIRD dev dataset. ACT-SQL Zhang et al. (2023a) cannot be evaluated on the BIRD dataset because it relies heavily on the Spider dataset. In addition, DIN-SQL Pourreza and Rafiei (2023) cannot be evaluated with Llama 3.1-8B-Instruct model because input prompts, provided in their official code, exceed the maximum token length of the model which is 8,192. As shown in the Table 10, existing methods exhibit similar performance to random demonstrations, while our method outperforms them, indicating its strong generalization performance.
Method EA VES
GPT-3.5-Turbo
DIN-SQL 45.1 52.4
DAIL-SQL 45.9 50.6
DCG-SQL 49.7 54.4
Deepseek-Coder-33B-Instruct
DIN-SQL 48.8 53.4
DAIL-SQL 49.9 52.6
DCG-SQL 53.2 56.0
Llama 3.1-8B-Instruct
DAIL-SQL 38.40 39.38
DCG-SQL 42.20 46.67
Deepseek-Coder-6.7B-Instruct
DIN-SQL 41.7 44.8
DAIL-SQL 40.7 43.8
DCG-SQL 44.8 47.2
Table 10: Performance on BIRD dataset with various LLMs. EA and VES indicate Execution Accuracy and Valid Efficiency Score, respectively.
Appendix B Analysis of Qualitative Results
To ensure a reliable qualitative analysis, we compare the generated SQL queries with those from existing approaches Gao et al. (2024); Zhang et al. (2023a) by using diverse cases of questions and DB schemas. Figure 3 and Figure 4 demonstrate the question, database, and selected demonstrations, which are used for SQL query generation. In the existing approaches, the given database contains all schema items, making the model choose irrelevant schema items for SQL query generation. Additionally, the selected demonstrations differ significantly in syntax from the target SQL query. Such difference cause the model to generate incorrect SQL queries since in-context learning makes LLMs follow the patterns in the examples Brown et al. (2020). In contrast, our method effectively includes relevant schema items and the selected demonstrations contain the highly similar to target SQL query. These indicate that our method successfully identifies the useful DB schema information and demonstrations for generating the target SQL.
Appendix C Prompt Details
In this section, we presents all the prompts of our methods. Our prompt is based on API DOCS format Rajkumar et al. (2022). Following Lin et al. (2020), we also provide potentially useful values by performing if there is match between question tokens and the values of each column.
C.1 Test Case Prompt
We demonstrate the prompt for a test case in Table 11.
C.2 Automated-CoT prompt
As shown in Table 12, we categorized the training samples based on the SQL into four cases, SIMPLE, JOIN, NESTED, and IUEN. We designed automated-CoT prompt beginning with "Let’s think step by step." except for SIMPLE cases. For JOIN, we make the model to generate intermediate step first, then generate final query Pourreza and Rafiei (2023). For nested and IEUN(INTERSECT, EXCEPT, UNION, NOT IN), we added a natural language phrase to describe briefly about the nested SQL query by converting WHERE clause and HAVING clause of the nested query with rule-based approach Tian et al. (2023).
Prompt
### Answer the question by SQLite SQL query only and with no explanation.
### SQLite SQL tables, with their properties:
# city (id, name, countrycode, district, population)
# sqlite_sequence (name,seq)
# country (code, name, continent, region, surfacearea, indepyear, population,
lifeexpectancy, gnp, gnpold, localname, governmentform, headofstate, capital, code2)
# countrylanguage (countrycode, language [Dutch, English], isofficial, percentage)
# Foreign Keys = [city.countrycode = country.code, countrylanguage.countrycode = country.code]
### Question: Which regions speak Dutch or English?
Table 11: Test Case Prompt Examples.
Prompt
Simple
### Answer the question by SQLite SQL query only and with no explanation.
### SQLite SQL tables, with their properties:
# head (age,…)
# Foreign Keys = […]
### Question: How many heads of the departments are older than 56 ?
### SQL: SELECT COUNT(*) FROM head WHERE age > 56 ;
JOIN
### Answer the question by SQLite SQL query only and with no explanation.
### SQLite SQL tables, with their properties:
# station (id, name, long, …)
# trip (duration, start_station_id, …)
# Foreign Keys = […]
### Question: For each station, return its longitude and the average duration of trips
that started from the station.
Let‘s think step by step. we need to join the tables ’station’ and ’trip’.
Create an intermediate representation, then use it to construct the query.
Intermediate representation: "FROM station AS T1 JOIN trip AS T2 ON T1.id = T2.start_station_id".
### SQL: SELECT T1.name , T1.long , AVG(T2.duration) FROM station AS T1 JOIN trip AS T2
ON T1.id = T2.start_station_id GROUP BY T2.start_station_id ;
NESTED
### Answer the question by SQLite SQL query only and with no explanation.
### SQLite SQL tables, with their properties:
# weather (date, min_dew_point_f, zip_code, ….)
# Foreign Keys = […]
### Question: On which day and in which zip code was the min dew point lower than
any day in zip code 94107?
Let‘s think step by step. we need a nested subquery for ’zip code of the weather is 94107’.
Nested subquery: "( SELECT min ( min_dew_point_f ) FROM weather WHERE zip_code = 94107 )".
With the nested subquery, we can get the final SQL query.
### SQL: SELECT date , zip_code FROM weather
WHERE min_dew_point_f < (SELECT MIN(min_dew_point_f) FROM weather
WHERE zip_code = 94107) ;
IUEN
### Answer the question by SQLite SQL query only and with no explanation.
### SQLite SQL tables, with their properties:
# city (status, population,…)
# Foreign Keys = […]
### Question: Show the status shared by cities with population bigger than 1500
and smaller than 500.
Let‘s think step by step. The question can be solved using the ’INTERSECT’ set operator
and two subqueries: one for ’population of the city is above 1500’ and the other for
’population of the city is less than 500’.
First subquery: SELECT Status FROM city WHERE Population > 1500
Second subquery: SELECT Status FROM city WHERE Population < 500
With the nested subquery, we can get the final SQL query.
### SQL: SELECT status FROM city WHERE population > 1500 INTERSECT
SELECT status FROM city WHERE population < 500 ;
Table 12: Automated-CoT for demonstrations.